# Bone Soups: A Seek-and-Soup Model Merging Approach for Controllable Multi-Objective Generation

### v1

@@ A.2.5 RQ5: How Does the Determination of Merging Coefficients Affect the Performance of Bone Soup?
A.2.6 RQ6: How Do Different Values of $\beta$ Impact the Performance of Bone Soup?
We conducted experiments on the different trade-offs in Long Form QA task and Reddit Summary task. In Figure 4 and Figure  13, we can see that by varying $\beta$, the resulting front consistently outperforms and dominates RS. This shows that the choice of $\beta$ does not significantly impact the performance of the front and a good $\beta$ only would further improve the upper bound of our method. This demonstrates the robustness in choosing the $\beta$ parameter, thereby underscoring both the lower bound performance and overall robustness of our method.
@@ A.2.6 RQ6: How Do Different Values of $\beta$ Impact the Performance of Bone Soup?
A.2.7 RQ7: What is the Impact of Extrapolation on Performance of Bone Soup?
@@
A.3 Experiments Setup Details
A.3.1 Baselines
@@
A.3.2 Long Form QA task
@@
A.3.3 Helpful Assistant Task
The Helpful Assistant task requires the model to generate appropriate responses based on a given user conversation history, ensuring the response is as helpful as possible while maintaining safety. We use the hh-rlhf dataset for training and evaluation. The hh-rlhf dataset consists of 160k prompts, responses, and corresponding human annotations, which is publicly available under the MIT License. Our use of the dataset is consistent with its intended use. Additionally, we use three open-source reward models following Yang et al. (2024) to assess the helpfulness, harmlessness, and humor of the responses generated by the model. The backbones of first two reward models are GPT-2 modelRadford et al. (2019). The two reward models were trained on the Anthropic/hh-rlhf dataset using pair-wise feedback. The harmless reward model achieves a test set accuracy of 0.73698 and the helpful reward model achieves an accuracy of 0.72621 on the test set. The humor reward model is a fine-tuned version of distilbert-base-uncased Sanh (2019) on a joke/no-joke dataset to detect humor. And there is no extra prompt for this task.
A.3.4 Reddit Summary Task
The Reddit Summary task Stiennon et al. (2020) focuses on summarizing Reddit posts, aiming to produce concise and coherent summaries that effectively capture the main content of the post. The dataset consists of Reddit threads, where the input comprises a post’s title and body text, and the output is a human-written summary, which is publicly available under the MIT License. Our use of the dataset is consistent with its intended use. And the prefix of the prompt is "Generate a one-sentence summary of this post." according to Yang et al. (2024). We use two open-source reward models Yang et al. (2024) to assess the quality of the summary generated from two different aspects.
A.3.5 Parameters Setting Details
@@
A.4 Evaluation Metrics Details
@@
A.5 Discussions
Merging Coefficients.
Solving the linear system may result in negative values, which could lead to negative interpolation of the backbone models. Our method Bone Soup of constructing non-orthogonal bases and then performing interpolation can thus be seen as a form of extrapolation. Unlike the naive merging method, where coefficients are constrained to be nonnegative, the presence of negative coefficients extends beyond the original model space. More importantly, the ability to interpret the negative coefficients selected offers the potential to improve performance in regions that would otherwise be inaccessible through standard interpolation techniques.

### v2

@@ A.3.5 RQ5: How Does the Determination of Merging Coefficients Affect the Performance of Bone Soup?
A.3.6 RQ6: How Do Different Values of $\beta$ Impact the Performance of Bone Soup?
We conducted experiments on the different trade-offs in Long Form QA task and Reddit Summary task. In Figure 4 and Figure 13, we can see that by varying $\beta$, the resulting front consistently outperforms and dominates RS. This shows that the choice of $\beta$ does not significantly impact the performance of the front and a good $\beta$ only would further improve the upper bound of our method. This demonstrates the robustness in choosing the $\beta$ parameter, thereby underscoring both the lower bound performance and overall robustness of our method.
@@ A.3.6 RQ6: How Do Different Values of $\beta$ Impact the Performance of Bone Soup?
A.3.7 RQ7: What is the Impact of Extrapolation on Performance of Bone Soup?
@@
A.3.8 Comparision with MODPO
Regarding MODPO Zhou et al. (2024), it is important to note that it is not an adaptive multi-objective generation method in the same sense as Bone Soup, MOD, RiC, or Rewarded Soup. MODPO requires re-training a model for each specific user preference, which is computationally expensive.
We also conduct additional experiments to include MODPO as a baseline for completeness. Since MODPO requires training for each preference, which incurs a huge overhead, we select values of $\alpha\in\{0.1,0.3,0.5,0.7,0.9\}$ for the user preference $(\alpha,1-\alpha)$. The experimental results are shown in Table 6 and Table 7:
As can be seen, Bone Soup consistently outperforms MODPO across various metrics. Although MODPO and Bone Soup do not share overlapping solutions, MODPO’s Pareto front is significantly smaller, covering a more limited solution space. This limitation is also reflected in its poorer hypervolume performance.
Based on our analysis, we suspect that the relatively weak performance of MODPO can be attributed to the lack of sufficient distinction of the reward in the dataset. As described in the original paper on MODPO, in addition to the margin reward, a second reward model is required to label the preference dataset, where the original "chosen" and "rejected" responses might have low distinction under the current reward model. This reduces the effectiveness of the DPO training process. On the other hand, Bone Soup, Rewarded Soup, and other similar methods utilize reinforcement learning to obtain diverse backbone models. The process of sampling responses in RL helps mitigate the issue of low distinction in rewards. This could be the reason why MODPO performs worse in our experiments.
Table 6: The comparison between Bone Soup and MODPO on Helpful vs Harmeless
Method HyperVolume Inner Product Controllability
Bone Soup 1.16 2.00 1.00
MODPO 0.93 -0.57 1.00
Table 7: The comparison between Bone Soup and MODPO on Faituful vs Preference1
Method HyperVolume Inner Product Controllability
Bone Soup 1.12 1.89 1.00
modpo 0.65 0.56 0.70
A.4 Experiments Setup Details
A.4.1 Baselines
@@
A.4.2 Long Form QA task
@@
A.4.3 Helpful Assistant Task
The Helpful Assistant task requires the model to generate appropriate responses based on a given user conversation history, ensuring the response is as helpful as possible while maintaining safety. We use the hh-rlhf dataset for training and evaluation. The hh-rlhf dataset consists of 160k prompts, responses, and corresponding human annotations, which is publicly available under the MIT License. Our use of the dataset is consistent with its intended use. Additionally, we use three open-source reward models following Yang et al. (2024)
to assess the helpfulness, harmlessness, and humor of the responses generated by the model. The backbones of first two reward models are GPT-2 modelRadford et al. (2019). The two reward models were trained on the Anthropic/hh-rlhf dataset using pair-wise feedback. The harmless reward model achieves a test set accuracy of 0.73698 and the helpful reward model achieves an accuracy of 0.72621 on the test set. The humor reward model is a fine-tuned version of distilbert-base-uncased Sanh (2019) on a joke/no-joke dataset to detect humor. And there is no extra prompt for this task.
A.4.4 Reddit Summary Task
The Reddit Summary task Stiennon et al. (2020) focuses on summarizing Reddit posts, aiming to produce concise and coherent summaries that effectively capture the main content of the post. The dataset consists of Reddit threads, where the input comprises a post’s title and body text, and the output is a human-written summary, which is publicly available under the MIT License. Our use of the dataset is consistent with its intended use. And the prefix of the prompt is "Generate a one-sentence summary of this post." according to Yang et al. (2024). We use two open-source reward models Yang et al. (2024)
to assess the quality of the summary generated from two different aspects.
A.4.5 Parameters Setting Details
@@
A.5 Evaluation Metrics Details
@@
A.6 Discussions
Computational Complexity Analysis The primary computational cost of our approach occurs during the training phase. Once training is completed, the inference stage can adapt to any user preference without additional overhead. Unlike MOD Shi et al. (2024) and MODPO Zhou et al. (2024), there is no extra cost during inference. The process of determining merging coefficients only requires solving a linear equation, which incurs no additional computational expense.
On the other hand, although we propose using a small-scale selection to choose the optimal $\beta\in\{0.8,0.7,0.6\}$, the robustness of the constructed matrix (see Appenidx A.3.3) means we are not overly dependent on $\beta$ and can get rids of this additional expense. In Appendix A.3.6, we discuss the impact of varying $\beta$ on Bone Soup, and the conclusion is that, regardless of the $\beta$ value, the resulting Pareto front consistently outperforms and dominates other baselines. Our insight is that including marginal reward could lead to better backbone construction. Therefore, the simplest approach is to select any $\beta$, which results in computational costs identical to those of standard Rewarded Soup, with no additional overhead. If one aims to achieve the optimal performance, let the computational cost of Rewarded Soup be $C$. Then, selecting optimal $\beta$ introduces an additional cost of $0.2\times C\times n$, where $n\leq 3$ is the number of possible choices for $\beta$ and as the small-scale selection trained the model with 20% total steps.
Negative Merging Coefficients.
Solving the linear system may result in negative values, which could lead to negative interpolation or said extrapolation Ilharco et al. (2022); Zheng et al. (2024) of the backbone models.
Previous works have discussed improving model performance through extrapolation techniques Zheng et al. (2024), or by deliberately weakening the initial SFT model to facilitate unlearning Ilharco et al. (2022). In these approaches, negative merging coefficients were determined through trial-and-error methods using a validation set.
However, in our approach, we avoid the cumbersome and somewhat unnatural trial-and-error process by directly solving linear equations to establish a clear mapping between the backbone models’ rewards and user preferences. This not only simplifies the process but can also be regarded as an interpretable extrapolation. Our method Bone Soup of constructing non-orthogonal bases and then performing interpolation can thus be seen as a form of extrapolation. Unlike the naive merging method, where coefficients are constrained to be nonnegative, the presence of negative coefficients extends beyond the original model space. More importantly, the ability to interpret the negative coefficients selected offers the potential to improve performance in regions that would otherwise be inaccessible through standard interpolation techniques.